Help solve for y.
6/y=7/y-1/y

Answer by stanbon(75887)

Help solve for y.
6/y=7/y-1/y
---
Multiply thru by y to get:
6 = 7-1
6 = 6
-----------------
Solution: y is any Real Number other than zero.
=================================================
Cheers,
Stan H.